# Interpretable LLM-based Table Question Answering

### v2

@@
Abstract.
Interpretability for Table Question Answering (Table QA) is critical, particularly in high-stakes industries like finance or healthcare. Although recent approaches using Large Language Models (LLMs) have significantly improved Table QA performance, their explanations for how the answers are generated are ambiguous. To fill this gap, we introduce Plan-of-SQLs (POS), an interpretable Table QA approach designed to improve users’ understanding of model decision-making.
Through qualitative and quantitative evaluations with human and LLM judges, we show that: First, POS is the highest-quality explanation method, helps human users understand model behaviors, and facilitates model prediction verification. Second, when evaluated on popular and standard Table QA datasets (TabFact, WikiTQ, and FetaQA), POS achieves QA accuracy that is competitive with or superior to existing methods, while also offering greater efficiency—requiring significantly fewer LLM calls and table database queries—and robust performance on large-sized tables. Finally, we observe high agreement (up to 90%) between LLMs and human users when making decisions based on the same explanations, suggesting that LLMs could serve as an effective proxy for humans in evaluating explanations. This finding enables faster, more affordable evaluation of AI explanations—possibly accelerating trustworthy AI research while maintaining reliable judgments on interpretability.
Reference Format:
, 19(1): XXX-XXX, 2026.
doi:XX.XX/XXX.XX *Work done during Summer internship at JP Morgan AI Research. This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. Copyright is held by the owner/author(s).
Artifact Availability:
The source code, data, and other artifacts have been made available at https://github.com/giangnguyen2412/table-qa-2025.
1. Introduction
@@
An estimated 38% of office tasks involve working with tables, often using Excel spreadsheets (Richardson, 2022), highlighting the need for advanced tools for tabular data analysis. LLM-powered Table QA models (those depicted in Figure 1) address this gap by enabling users to quickly extract insights or answer questions for tables, making them invaluable in various industries. For example, financial analysts leverage these models to predict trends in tabular market data (Lo and Ross, 2024). Similarly, medical professionals use them to analyze tabular medical records of patients, facilitating accurate and timely treatment decisions (Bardhan et al., 2022).
@@
However, the value of these systems comes with high risks. Errors in financial decision making have led to catastrophic outcomes, such as the billion-dollar loss Citigroup faced in 2022 (Jane, 2024). In healthcare, the stakes are even fatal, considering a model that misjudges a young man’s health by overlooking his family history, resulting in his death from cardiac arrest just weeks later (Stanford HAI, 2024). These examples highlight the pressing need for interpretability in Table QA to ensure safe and accountable decisions (Fang et al., 2024).
@@
Despite the importance, interpretability remains an underexplored dimension in Table QA literature. Recent approaches have significantly increased accuracy and often present themselves as interpretable solutions (Ye et al., 2023; Cheng et al., 2023; Wang et al., 2024), but this supposed interpretability remains unsubstantiated by empirical evidence. Anecdotally, when visualizing their explanations in Figure 1c, we observe that model behaviors may not be understandable to users (e.g., a user may wonder why certain rows are selected in f_select_row() or how simple_query() generates the answer).
@@
Figure 1. (a) End-to-End: relies entirely on an LLM to answer the question directly, leaving no room for users to understand the prediction. (b) Text-to-SQL: gGenerates an SQL command to solve the query, requiring domain expertise to understand and becoming unintelligible when the query becomes complex. (c) Chain-of-Table or CoTable: performs planning with abstract functions and executes sequentially to arrive at the final answer. However, function arguments are not justified, and the final answer depends on the LLM’s opaque reasoning. (d) Plan-of-SQLs or POS (Ours): plans in natural language, making each step simple and understandable. Each step is then converted into an SQL command, sequentially transforming the input table end-to-end to produce the final answer. We provide a publicly accessible interface to compare all explanations for the same query.
To bridge this gap, we propose Plan-of-SQLs or POS, an interpretable Table QA approach that decomposes the original query into a sequence of atomic SQL operations—each individually simple, self-contained, and limited to a single straightforward table transformation. This has two main benefits. First, POS prevents the model from arbitrarily selecting irrelevant data (Wang et al., 2024) (row 3 in Figure 1c) as seen in hybrid approaches via explicit condition-based filtering in SQL (e.g., step 2 in Figure 1d). Second, POS directly addresses the opacity in the answer-generation step ( Figure 1a & c) by a trivial SQL command execution (step 3 in Figure 1d).
To rigorously investigate the actual interpretability of POS, we perform both qualitative and quantitative evaluations, comparing POS with baselines that are considered interpretable, including Text-to-SQL (Rajkumar et al., 2022), DATER (Ye et al., 2023), and Chain-of-Table (CoTable) (Wang et al., 2024). In three evaluation benchmarks (quality preference ranking (Ramaswamy et al., 2023; Yang et al., 2024), forward behavior simulation (Doshi-Velez and Kim, 2017; Hase and Bansal, 2020; Chen et al., 2022; Mills et al., 2023), and model prediction verification (Nguyen et al., 2021b; Taesiri et al., 2022)) with human and LLM judges, POS consistently receives the highest ratings and helps users understand and verify model decisions substantially better.
When tested against popular and standard Table QA datasets such as TabFact (Chen et al., 2020), WikiTQ (Pasupat and Liang, 2015), and FetaQA (Nan et al., 2022), POS achieves competitive or superior accuracy compared to existing methods, uses much fewer LLM calls and database queries, and shows robust performance on large-sized tables.
In summary, the key contributions of our paper are:
  • We are the first in the LLM-based Table QA literature to study the effectiveness of explanation methods and to propose a novel explanation (POS) that facilitates users’ understanding of model decision-making. (details in  Section 3).
@@
  • Compared to existing explainable Table QA methods, POS is more robust when handling large tables and more efficient in its use of LLM calls and database queries, while still achieving competitive or superior Table QA performance (details in  Section 4).
@@
  • Our experiments reveal the high agreement between humans and LLMs in evaluating Table QA explanations (up to 90%), suggesting that LLMs can serve as effective and scalable proxy users for humans in Explainable AI (XAI) evaluation (details in  Section 4.1.3).
2. Related Work
2.1. Decomposing Input Queries in Table QA
@@
LLM-based Table QA models have improved performance by decomposing complex input queries into smaller problems (Ye et al., 2023; Nahid and Rafiei, 2024b; Zhao et al., 2024) or by employing step-by-step reasoning (Wang et al., 2024; Wu and Feng, 2024; Abhyankar et al., 2024). However, these approaches often rely on highly complex table transformations—for instance, selecting a query-relevant subtable from a large input table via the opaque reasoning of LLMs (Ye et al., 2023; Nahid and Rafiei, 2024b; Wu and Feng, 2024; Abhyankar et al., 2024)—which can lead to errors in retrieving the necessary data (see Appendix E for examples of hallucination in sub-table selections). Often, the resulting sub-tables contain entries that are either irrelevant or logically incorrect due to the long-context challenges inherent in black-box LLM reasoning.
@@
In contrast, POS constrains every transformation of the table to an atomic SQL operation, that is, a single clause with at most one condition and one variable (e.g. “Select rows where opponents = 0”). This design has two main advantages. First, high accuracy and interpretability: Because each step is minimal, there is less room for the LLM to make logical errors; at the same time, the transformation is transparent and easy to follow for users. Second, fine-grained attribution: Executing atomic SQLs enables direct identification of which cells are used per step, providing a detailed account of how the final answer is derived from the input.
2.2. Program-based Table Transformations
Using programming languages such as SQL (Nahid and Rafiei, 2024b; Ye et al., 2023) or Python (Cheng et al., 2023; Chen et al., 2020) for table transformations offers two key benefits. First, these languages enable rule-based transformations that provide great traceability of input features being used, thereby enhancing interpretability compared to the black-box operations of LLMs (see Figure 1a and c). Second, these programming languages facilitate efficient handling of complex table operations, making them more reliable and cost-effective than LLM-based methods. In contrast, LLM-based approaches are inefficient and prone to errors when processing entire large tables due to the intrinsic limitations of managing long contexts in black-box LLM reasoning (Chen, 2023; Wang et al., 2024).
@@ 2.2. Program-based Table Transformations
Our work contributes to a growing body of literature that harnesses program-based table transformations for Table QA but uses SQL commands exclusively. To our knowledge, only two methods in the Table QA literature, LPA (Chen et al., 2020) (using Python-Pandas) and Text-to-SQL (Rajkumar et al., 2022), handle queries end-to-end through program-based operations. However, since Text-to-SQL generates a single SQL command for a query ( Figure 1b), it requires a highly powerful Text-to-SQL converter and often generates error-prone SQL commands (Shi et al., 2020). Meanwhile, LPA constructs a single Python-Pandas program to represent the entire query, which can lead to complexity and potential errors due to the challenges of synthesizing accurate programs in one forward pass. In contrast, our approach breaks down queries into multiple atomic steps expressed in natural language, which are then easily translated into SQL. This eliminates the need for powerful Text-to-SQL models while achieving superior accuracy (see Table 6).
2.3. Evaluating Interpretability
@@
Interpretability is a critical yet underexplored aspect of the LLM-based Table QA literature. We contribute to a substantial body of work that evaluates machine explanations with human users (Adebayo et al., 2020; Nguyen et al., 2021b; Kim et al., 2022; Taesiri et al., 2022; Colin et al., 2022; Steyvers and Kumar, 2023; Chen et al., 2023a; Nguyen et al., 2024a, b; Zhao et al., 2024; Zhang et al., 2025). Although TAPERA (Zhao et al., 2024) evaluated explanations in LLM-based Table QA by having participants subjectively rate their faithfulness and comprehensiveness on a 5-point Likert scale, these two metrics primarily serve as proxies for explanation quality rather than directly measuring their practical effectiveness for human users. In addition, POS explanations explicitly represent the model’s reasoning process through offline SQL executions, which guarantees absolute (i.e., 100%) faithfulness and comprehensiveness. Therefore, we have opted not to evaluate these two metrics.
3. POS: Interpretable Table QA
@@
In Table Question Answering (Table QA), each instance is represented as a triplet $(T,Q,A)$, where $T$ is a table, $Q$ is a query about the table, and $A$ is the answer. The task is to predict the answer $A$ given the query $Q$ and the table $T$. POS decomposes $Q$ into sub-queries $\rightarrow$ converts each sub-query into an SQL command $\rightarrow$ applies these commands sequentially to $T$ to arrive at $A$.
Atomicity in Table QA.
An atomic step is defined as a simple, minimal operation that can be easily converted into a single SQL command. Specifically, each converted SQL command must: (i) contain at most one condition (e.g., in WHERE clause); and (ii) involve at most one variable (e.g., column name) within that condition. With step atomicity, we ensure that the converted SQL commands remain straightforward and free from errors during translation and execution. Additionally, atomic steps facilitate clear communication of the model’s reasoning process to users.
Figure 2. Illustration of Plan-of-SQLs (POS). 1 The NL Atomic Planner takes $(T,Q)$ as input and generates a step-by-step plan in natural language . 2 Step-to-SQL takes $(T,S1)$ as input and converts $S1$ to SQL to sort the table (transform $T\rightarrow T1$). 3 Step-to-SQL takes $(T1,S2)$ as input and converts $S2$ to SQL to select relevant rows (transform $T1\rightarrow T2$). 4 Step-to-SQL takes $(T2,S3)$ as input and converts $S3$ to SQL to return the final answer. An interactive interface of POS is publicly available.
3.1. Planning in Natural Language
We propose performing planning in natural language, a strategy that leverages LLMs’ inherent strengths—since they are primarily trained on natural text (Huang et al., 2022). This also makes the planning process more interpretable for humans, as each step is expressed in human-understandable terms rather than abstract function calls (e.g., simple_query), whose motivations and arguments are ambiguous (Wang et al., 2024).
In Figure 21, the Natural-Language (NL) Atomic Planner takes $(T,Q)$ as input and converts $Q$ into a plan of atomic steps. POS uses the following prompt to create the plan, incorporating atomicity through both in-context planning examples and instructions:

### v3

@@
Abstract
Interpretability in Table Question Answering (Table QA) is critical, especially in high-stakes domains like finance and healthcare. While recent Table QA approaches based on Large Language Models (LLMs) achieve high accuracy, they often produce ambiguous explanations of how answers are derived. We propose Plan-of-SQLs (POS), a new Table QA method that makes the model’s decision-making process interpretable. POS decomposes a question into a sequence of atomic steps, each directly translated into an executable SQL command on the table, thereby ensuring that every intermediate result is transparent. Through extensive experiments, we show that: First, POS generates the highest-quality explanations among compared methods, which markedly improves the users’ ability to simulate and verify the model’s decisions. Second, when evaluated on standard Table QA benchmarks (TabFact, WikiTQ, and FeTaQA), POS achieves QA accuracy that is competitive to existing methods, while also offering greater efficiency—requiring significantly fewer LLM calls and table database queries (up to 25$\times$ fewer)—and more robust performance on large-sized tables. Finally, we observe high agreement (up to 90.59% in forward simulation) between LLMs and human users when making decisions based on the same explanations, suggesting that LLMs could serve as an effective proxy for humans in evaluating Table QA explanations. Code and data available at: https://github.com/anguyen8/pos
1 Introduction
@@
An estimated 38% of office tasks involve working with tables, often using Excel (Richardson, 2022), highlighting the need for advanced tools for tabular data analysis. LLM-powered Table QA models (those in Fig. 1) address this gap by enabling users to quickly extract insights or answer questions for tables, making them invaluable in various industries. For example, financial analysts leverage these models to predict trends from tabular market data (Lo & Ross, 2024). Similarly, medical professionals use them to analyze tabular medical records of patients, facilitating accurate and timely treatment decisions (Bardhan et al., 2022).
@@
However, the value of these systems comes with high risks. Errors in financial decision making have led to catastrophic outcomes, such as the billion-dollar loss Citigroup faced in 2022 (Jane, 2024). In healthcare, the stakes are even fatal, considering a model that misjudges a man’s health by overlooking his family history, resulting in his death from cardiac arrest weeks later (Stanford HAI, 2024). These examples underscore the pressing need for interpretability in Table QA to ensure safe and accountable use of AI (Fang et al., 2024).
@@
Despite its importance, interpretability remains an underexplored dimension in Table QA literature. Recent approaches have significantly increased accuracy and often present themselves as interpretable solutions (Ye et al., 2023; Cheng et al., 2023; Wang et al., 2024), but this interpretability is unsubstantiated by empirical evidence. In practice, the explanations provided by these models can be unclear. For instance, as shown in Fig. 1c, a user cannot discern why certain rows were selected by a function f_select_row() or how an operation like simple_query() produced the final answer. In other words, current Table QA methods do not adequately explain their reasoning to users.
@@
Figure 1: (a) End-to-End: relies entirely on an LLM to answer the question directly, leaving no room for users to understand the prediction. (b) Text-to-SQL: generates an SQL command to solve the question, requiring domain expertise to understand and becoming unintelligible when the question becomes complex. (c) Chain-of-Table or CoTable: performs planning with abstract functions and executes sequentially to arrive at the final answer. However, function arguments are not justified, and the final answer depends on the LLM’s opaque reasoning. (d) Plan-of-SQLs or POS (Ours): plans in natural language, making each step simple and understandable. Each step is then converted into an SQL command, sequentially transforming the input table end-to-end to produce the final answer. We provide a public interface to compare explanations.
To address this gap, we propose Plan-of-SQLs (or POS), an LLM-based Table QA approach that places interpretability at its core. POS decomposes each question into a sequence of atomic steps, where each step is a simple sub-problem that can be translated into a corresponding SQL command and executed on the table. By design, each transformation is self-contained and limited in scope, forcing the model’s decision-making to be broken down into transparent and verifiable steps. This has two key benefits. First, by requiring simple step-by-step table transformations via SQL, we avoid the model arbitrarily pulling in irrelevant data for answering. For example, existing methods that select a large subtable in one shot (Ye et al., 2023; Wang et al., 2024) often include spurious entries due to the black-box LLM reasoning (see the irrelevant selection of row3 in Fig. 1c). By contrast, POS uses programmatic conditions (e.g., “WHERE opponents = 0”) to ensure only relevant entries are chosen (Step 2 of Fig. 1d). Second, POS largely avoids the opaque answer-generation step commonly found in LLM-based Table QA literature (Ye et al., 2023; Wang et al., 2024; Nahid & Rafiei, 2024b; Zhao et al., 2024) (illustrated in Fig. 1a,c). Rather than sending the entire table to an LLM, POS produces its final answer via a deterministic SQL query execution (Step 3 of Fig. 1d), making it entirely clear how the answer is derived from the data.
We thoroughly evaluate the interpretability and performance of POS. We compare against existing Table QA methods that provide explanations, including Text-to-SQL (Rajkumar et al., 2022), DATER (Ye et al., 2023), Chain-of-Table (CoTable)(Wang et al., 2024), and self-explaining (Madsen et al., 2024). Our interpretability evaluation spans three benchmarks—(1) explanation quality preference ranking (Ramaswamy et al., 2023; Yang et al., 2024), (2) forward simulation of the model’s behavior (Doshi-Velez & Kim, 2017; Hase & Bansal, 2020; Chen et al., 2022; Mills et al., 2023), and (3) model prediction verification (Nguyen et al., 2021; Taesiri et al., 2022)—using both human participants and LLM-based judges. In all settings, POS usually yields the best results, showing that its explanations help users (human or AI) simulate and verify the model output far more effectively than the existing interpretable competitors. Furthermore, when tested on standard datasets (TabFact (Chen et al., 2020), WikiTQ (Pasupat & Liang, 2015), and FeTaQA (Nan et al., 2022)), POS achieves accuracy on par (within 1–2 points) with existing methods, while requiring drastically fewer LLM calls and database queries. Notably, POS scales robustly to large tables, where current methods Cheng et al. (2023); Ye et al. (2023); Wang et al. (2024) often struggle. Finally, we observe a high agreement between human evaluators and LLM-based evaluators (up to 90%) in forward simulation, suggesting that LLM-based proxies can reliably stand in for human judgment. In summary, our contributions are as follows.
  • We introduce POS, a new Table QA method that is designed for interpretability. We carry out a thorough study of explanation effectiveness, and show that our explanations substantially improve users’ understanding of the model’s decision-making over existing interpretable methods (see Sec. 4.1.3).
@@
  • Compared to existing Table QA approaches, POS is more robust on large tables and significantly more efficient in its use of LLM calls and database queries (25$\times$ fewer LLM calls than DATER, and 25$\times$ fewer database queries than Binder), while maintaining competitive QA accuracy (see Sec. 4.2).
@@
  • Our experiments reveal high agreement between human users and LLM judges in evaluating Table QA explanations (up to 90.59% agreement in forward simulation). This suggests that LLMs could effectively proxy for human evaluators in evaluating Table QA explanations (see Sec. 4.1.5).
2 Related Work
2.1 Atomic Table Transformations for Table QA
@@
LLM-based Table QA models have improved performance by decomposing complex input queries into smaller problems (Ye et al., 2023; Nahid & Rafiei, 2024b; Zhao et al., 2024) or by employing step-by-step reasoning (Wang et al., 2024; Wu & Feng, 2024; Abhyankar et al., 2024). However, these approaches often rely on highly complex table transformations—for instance, selecting a question-relevant subtable from a large input table via the opaque reasoning of LLMs (Ye et al., 2023; Nahid & Rafiei, 2024b; Wu & Feng, 2024; Abhyankar et al., 2024)—which can lead to errors in retrieving the correct data (see details in Appendix F). Often, the resulting sub-tables contain entries that are either irrelevant or logically incorrect with respect to the input question, due to the long-context challenges inherent in LLM reasoning.
@@
In contrast, POS constrains a transformation of the table to an atomic SQL operation, that is, a single clause with at most one condition and one variable (e.g., “Select rows where opponents = 0”). This design has two main advantages. First, it improves accuracy and comprehensibility: because each step is minimal and focused, there is less room for LLMs to make mistakes, and the operation is easy for users to digest. Second, it enables fine-grained attribution: by executing each step with an atomic SQL query, we can pinpoint exactly which cells were used at that step, yielding a detailed trace of how the final answer is derived.
2.2 Program-based Table Transformations
Using programming languages such as SQL (Nahid & Rafiei, 2024b; Ye et al., 2023; Stoisser et al., 2025) or Python (Cheng et al., 2023; Chen et al., 2020) to manipulate tables is preferable for two reasons. First, these languages perform rule-based operations with explicit references to table cells, offering much greater traceability than the implicit, black-box transformations of LLMs (see contrastive examples between Fig. 1a & d). Second, programmatic transformations can handle large or complex table operations more reliably and efficiently, since they do not suffer from the context length and inconsistency issues that LLM-based methods encounter when processing entire tables (Chen, 2023; Wang et al., 2024; Nahid & Rafiei, 2024b).
@@ 2.2 Program-based Table Transformations
POS builds on this line of work by leveraging program-based transformations—specifically, SQL—to solve Table QA. Notably, POS uses SQL exclusively for executing reasoning steps. To our knowledge, only two methods in the Table QA literature, LPA (Chen et al., 2020) (using Python-Pandas) and Text-to-SQL (Rajkumar et al., 2022), handle queries end-to-end through program-based operations. However, since Text-to-SQL generates a single complex SQL command for an input question (Fig. 1b), it requires a highly powerful Text-to-SQL model and often generates error-prone commands (Shi et al., 2020). Similarly, LPA’s one-pass program synthesis can be brittle, as generating a correct multi-step program in one go is inherently challenging (Chen et al., 2020). In contrast, POS breaks the problem into multiple simpler SQL queries corresponding to atomic sub-steps initially expressed in natural language. This stepwise use of SQL removes the need for an advanced program generator and, as our results will show in Sec. 4, leads to higher overall accuracy and interpretability than Text-to-SQL or LPA.
2.3 Evaluating Interpretability
@@
Interpretability is a critical aspect of AI models, and there is a rich body of work on evaluating explanations with human users (Adebayo et al., 2020; Nguyen et al., 2021; Kim et al., 2022; Taesiri et al., 2022; Colin et al., 2022; Steyvers & Kumar, 2023; Chen et al., 2023a; Nguyen et al., 2024a; b; Zhang et al., 2025). In the context of LLM-based Table QA, TAPERA (Zhao et al., 2024) had users subjectively rate the faithfulness and comprehensiveness of explanations on a Likert scale. However, such ratings serve only as proxies for explanation quality and do not directly measure the explanations’ utility to users. POS guarantees 100% faithfulness (the explanation steps exactly produce the answer) and comprehensiveness (no reasoning step is hidden) because its explanations explicitly represent the model’s reasoning through actual executed SQL operations. Therefore, in our experiments, we prioritize direct evaluation of explanations’ effectiveness—such as how well users can simulate or verify the model’s behaviors—over faithfulness or completeness metrics.
3 POS: Interpretable Table QA with Atomic Table Transformations
@@
In Table Question Answering (Table QA), each sample is represented as a triplet $(T,Q,A)$, where $T$ is a table, $Q$ is a question about the table, and $A$ is the answer. The task is to predict the answer $A$ given the question $Q$ and the table $T$. POS decomposes $Q$ into steps $\rightarrow$ converts each step into an SQL command $\rightarrow$ applies these commands sequentially to $T$ to arrive at $A$.
Atomicity in Table QA Reasoning.
We define an atomic step as a simple, minimal natural-language table operation that can be reliably translated into an SQL query. Specifically, each operation is restricted to: (i) at most one condition (e.g., =), and (ii) at most one column or variable in that condition (e.g., opponents). Enforcing this atomicity keeps each step’s translation and execution reliable and lowers the chance of errors. It also makes the model’s reasoning more interpretable: because each step addresses only a small part of the problem, a human can easily understand the purpose and effect of that step.
Figure 2: Illustration of Plan-of-SQLs (POS). 1 The Natural Language (NL) Atomic Planner takes $(T,Q)$ as input and generates a step-by-step plan in plain language. 2 Step-to-SQL then takes $(T,S_{1})$ and converts the first step $S_{1}$ into an SQL query, which is executed on $T$ to produce an intermediate table $T_{1}$. 3 Step-to-SQL takes $(T_{1},S_{2})$ to produce and execute the next SQL query, yielding $T_{2}$. 4 The final Step-to-SQL uses $(T_{2},S_{3})$ to generate an SQL query that returns the final answer. We provide an interactive demo interface of POS.
3.1 Planning in Natural Language
Rather than planning in a space of abstract functions for Table QA like Chen et al. (2020) and Wang et al. (2024), we perform the reasoning-step generation in natural language. Leveraging an LLM to plan in natural language takes advantage of the model’s strong priors from its language training (Huang et al., 2022). It also makes the plan inherently understandable: each step is described in natural language, which is far more interpretable than, say, a sequence of function names (e.g., simple_query()) whose purpose is not immediately clear (Wang et al., 2024). In Fig. 21, the NL Atomic Planner takes the input $(T,Q)$ and produces a plan of atomic steps in natural language that outlines how to derive the answer $A$. POS prompts an LLM with examples and explicit instructions to incorporate the atomic decomposition. Specifically, we use a prompt of this form: